please suggest solution:    The flat spiral has the smaller diameter = 5mm & major diameter = 15.5mm, & has 7.5 coils. It is required to find the total length of the wire to form a spiral  of the above description. Pl. help.

Grade:

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
We need to find the circumference of the flat spiral coil to calculate the length of wire.
Inner Diameter (d) = 5 mm
Outer Diameter (D) = 15.5 mm
Number of turns (N) = 7.5
Circumference’L’:
L = \pi N (\frac{D+d}{2})
L = \pi \times 7.5 (\frac{15.5+5}{2})
L = \pi \times 7.5 \times 10.25
L = 241.4 mm
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty